OpenStudy (anonymous):

How can this multiplication problem be written using only improper fractions? five times two and one-eighth equals what value? five over one times sixteen over eight equals what value? five over one times eight over sixteen equals what value? five over one times seventeen over eight equals what value? five over one times eighteen over seven equals what value?

OpenStudy (anonymous):

An improper fraction is a fraction where the numerator is bigger than the denominator. I'll do the first problem and lets see if you can go from there. \[2\frac{ 1 }{ 8 } = \frac{ 17 }{8 }\] I know this because 2 is the same as 16/8 and then add 1/8 to get 17/8. Just multiply the numerator by 5 and you get: \[\frac{ 85 }{ 8 }\]

OpenStudy (jdoe0001):

\(\Large \textit{proper} \qquad a\cfrac{b}{c}\qquad \textit{improper}\qquad \cfrac{a\times c+b}{c}\)